OpenStudy (anonymous):

The length of a rectangular room is 6 feet longer than twice the width. If the room's perimeter is 144 feet, what are the room's dimensions? A) Width = 27 ft; length = 60 ft B) Width = 33 ft; length = 39 ft C) Width = 22 ft; length = 50 ft D) Width = 44 ft; length = 100 ft

OpenStudy (anonymous):



OpenStudy (anonymous):

44ft by 100

OpenStudy (anonymous):

Perimeter is adding all the sides together, so 2w+6+2w+6+w+w=144 Solve for w

OpenStudy (anonymous):

6w+12=144

OpenStudy (anonymous):

6w=144-12

OpenStudy (anonymous):

22

OpenStudy (anonymous):

Right, thats the width. Since the length is 2w+6 length=2(22)+6 length=50 feet

OpenStudy (anonymous):

thank you i need to check where i went wrong.